### v1

# Generate-on-Graph: Treat LLM as both Agent and KG in Incomplete Knowledge Graph Question Answering

Abstract
To address the issue of insufficient knowledge and the tendency to generate hallucination in Large Language Models (LLMs), numerous studies have endeavored to integrate LLMs with Knowledge Graphs (KGs). However, all these methods are evaluated on conventional Knowledge Graph Question Answering (KGQA) with complete KGs, where the factual triples involved in each question are entirely covered by the given KG. In this situation, LLM mainly acts as an agent to find answer entities by exploring the KG, rather than effectively integrating internal and external knowledge sources. However, in real-world scenarios, KGs are often incomplete to cover all the knowledge required to answer questions. To simulate real-world scenarios and evaluate the ability of LLMs to integrate internal and external knowledge, in this paper, we propose leveraging LLMs for QA under Incomplete Knowledge Graph (IKGQA), where the given KG doesn’t include all the factual triples involved in each question. To handle IKGQA, we propose a training-free method called Generate-on-Graph (GoG) that can generate new factual triples while exploring on KGs. Specifically, we propose a selecting-generating-answering framework, which not only treat the LLM as an agent to explore on KGs, but also treat it as a KG to generate new facts based on the explored subgraph and its inherent knowledge. Experimental results on two datasets demonstrate that our GoG can solve IKGQA to a certain extent, while almost all previous methods cannot perform well on IKGQA.
1 Introduction
Figure 1: An example of Incomplete Knowledge Graph Question Answering (IKGQA), where yellow and red nodes represent topic and answer entity, respectively. Red dash line represents this triple (Cupertino, timezone, Pacific Standard Time) is missing in the KG.
Large Language Models (LLMs) Brown et al. (2020); Bang et al. (2023) have made great success in various natural language processing (NLP) tasks. Benefiting from extensive model parameters and vast amounts of pre-training corpus, LLMs can solve complex reasoning tasks through In-Context Learning (ICL) Dong et al. (2023), without fine-tuning for specific tasks.
However, LLMs are still suffer from insufficient knowledge and the tendency to generate hallucinationHuang et al. (2023); Li et al. (2023a). To mitigate this issue, many methods that incorporate LLMs with Knowledge Graphs (KGs) Ji et al. (2021) has been proposed Pan et al. (2023), where KGs provide accurate and abundant factual knowledge in triple format while LLMs provide strong natural language processing ability. These works can be roughly divided into two categories: (1) Semantic Parsing (SP) methods Li et al. (2023c); Nie et al. (2023), which use LLMs to generate logical forms with ICL, and then obtain answers by executing these logical queries on KGs. (2) Retrieval Augmented (RA) methods Li et al. (2023d), which retrieve information related to the question from KGs as external knowledge to help LLMs to obtain the answers. The retrieval format can be triples Li et al. (2023d), paths Sun et al. (2023); Luo et al. (2023b) and subgraphs Wang et al. (2023a).
Semantic parsing based methods exclusively treat LLMs as parser, not only do they ignore LLMs’ inherent knowledge and reasoning ability, but they also depend heavily on KGs’ quality and completeness Sun et al. (2023). Therefore, more attention is paid to retrieval augmented methods.
Figure 2: The performance of ToG Sun et al. (2023) and our GoG under complete/incomplete KG. Dash line indicates the performance of LLM with CoT Wei et al. (2023) prompting (without KG).
Although these retrieval augmented methods claim to solve the drawbacks of semantic parsing methods and obtain good performance on conventional Knowledge Graph Question Answering (KGQA) Yih et al. (2016a), it is still hard to determine if they effectively integrate knowledge from KGs and LLMs, and how much reasoning ability is utilized. One crucial reason is that, in conventional KGQA tasks, the factual triplets involved in each question are entirely covered by the KG. Therefore, in this scenario, LLMs still play the role of parser which only needs to identify the relationship path starting from the topic entity to the answer entity. For example, for the question "What is the timezone of the area where Apple headquarters is located?" in Figure 1, the LLM only needs to start from Apple headquarters, sequentially choose the relations located_in and timezone to find the answer. That means LLMs only need to ground the relationship appearing in the question to the specific relation in the KG to reach the answer entity Pacific Standard Time. Therefore, in conventional KGQA, LLM mainly acts as an agent to find answer entities by exploring the KG, rather than effectively integrating internal and external knowledge sources.
However, in real-world scenarios, KGs are often incomplete to cover all the knowledge required to answer questions. Besides, LLMs contain rich knowledge content and have powerful reasoning ability. Therefore, to evaluate the ability of LLMs to integrate internal and external knowledge, in this paper, we propose leveraging LLMs for QA under incomplete KG (IKGQA). The distinction between IKGQA and conventional KGQA is that IKGQA does not encompass all the factual triplets relevant to each question, rendering the KG in IKGQA incomplete. For example, as shown in Figure 1, for answering the same question "What is the timezone of the area where Apple headquarters is located?", the key triple (Cupertino, timezone, Pacific Standard Time) does not exist in the KG. This means that even if the correct SPARQL query is generated, it may not retrieve the final answer  [^1] SP methods do not parse ”timezone” into two hop path ”located_in -¿ timezone” because in the training set, ”timezone” only corresponds to one hop path ”timezone”, more details can be found in Appendix A.. Compared to KGQA, IKGQA holds greater research significance for the following reasons: (1) it is closer to real-world scenarios where the given KG is incomplete to answer users’ questions. (2) it evaluates LLMs’ reasoning ability and its capability to integrate inherent and external knowledge.
As shown in Figure 2, although previous methods have achieved outstanding performance in KGQA, their performance has significantly declined in IKGQA, even is worse than that without KG. This discrepancy suggests that these methods might not fully integrate the external and inherent knowledge of LLMs as professed.
To address the challenges in IKGQA, we propose a novel method called Generate-on-graph (GoG), which adopts a selecting-generating-answering framework consisting of three main steps: (1) Selecting, where LLMs select the relations most relevant to the current question and expanding the subgraph using these relations. (2) Generating, where LLMs use its inherent knowledge and reasoning abilities to generate new factual triples based on the explored subgraph. For example, if the KG already contains the triples (Cupertino, located_in, California), (California, timezone, Pacific Standard Time), LLMs can infer new triple (Cupertino, timezone, Pacific Standard Time). (3) Answering, where LLMs try to answer the question based on the retrieval and generated knowledge. If the information is still insufficient, the process is repeated by going back to Steps 1 and 2 until the LLMs can answer the question. The codes and data are available at https://github.com/YaooXu/GoG.
The main contributions of this paper can be summarized as follows:
  1. 1.
    We propose leveraging LLMs for QA under incomplete KG (IKGQA), which is closer to real-world scenarios and can evaluate LLMs’ ability better, and construct relevant datasets.
  2. 2.
    We propose a method called Generate-on-Graph (GoG), which uses the selecting-generating-answering framework, to address IKGQA.
  3. 3.
    Experimental results on two datasets show the superiority of GoG, and demonstrate that an incomplete KG can still help LLMs answer complex questions by providing related structured information, even without directly providing the answers.
Figure 3: Comparison of four methods in solving IKGQA: (a) Standard CoT prompting (closed-book, without KG), (b) Semantic parsing based method (e.g., KB-BINDER Li et al. (2023c)), (c) Path retrieval method (e.g., ToG Sun et al. (2023)), (d) The proposed GoG with selecting-generating-answering framework.
2 Related Work
Question Answering Under Incomplete KG. Some previous works Saxena et al. (2020); Zan et al. (2022) attempt to train KG embeddings to predict answers by similarity scores under incomplete KG. Compared to these previous KGE-based works, we propose leveraging LLMs for QA under incomplete KG to study whether LLMs can integrate internal and external knowledge well.
Unifying KGs and LLMs for KGQA. Various methods have been proposed to unify KGs and LLMs to solve KGQA, these methods can be classified into two categories: Semantic Parsing (SP) methods Li et al. (2023c); Nie et al. (2023) and Retrieval Augmented (RA) methods Luo et al. (2023b); Sun et al. (2023). SP methods transform the question into a structural query using LLMs. These queries can then be executed by a KG engine to derive answers based on KGs Sun et al. (2020). KB-BINDER Li et al. (2023c) generates the drafts as preliminary logical forms first, and then binds the drafts to the executable ones with entity and relation binders. However, the effectiveness of these methods relies heavily on the quality of the generated queries and the completeness of KGs. RA methods retrieve related information from the KG to improve the reasoning performance Li et al. (2023b). ToG Sun et al. (2023) treats the LLM as an agent to interactively explore relation paths step-by-step on KGs and perform reasoning based on the retrieved paths. RoG Luo et al. (2023b) first generates relation paths as faithful plans, and then use them to retrieve valid reasoning paths from the KGs for LLMs to reason. Our GoG belongs to retrieval augmented methods, we also utilize the knowledge modeling ability of LLMs, as well as the semantic parsing ability and reasoning ability.
LLM reasoning with Prompting. Many works have been proposed to elicit the reasoning ability of LLMs to solve complex tasks through prompting Wei et al. (2023); Khot et al. (2023). Complex CoT Fu et al. (2023) creates and refine rationale examples with more reasoning steps to elicit better reasoning in LLMs. Self-Consistency Wang et al. (2023b) fully explores various ways of reasoning to improve their performance on reasoning tasks. DecomP Khot et al. (2023) solves complex tasks by instead decomposing them into simpler sub-tasks and delegating these to sub-task specific LLMs. ReAct Yao et al. (2023) treats LLMs as agents that interact with the environment and make decisions to retrieve information from external source. GoG can be viewed as a fusion of ReAct and DecomP, thereby enabling a more comprehensive utilization of the diverse capabilities inherent in LLMs for addressing complex questions.
3 Preliminary
In this section, we firstly introduce Knowledge Graph (KGs). Then, we use symbols of KGs to define relation path and Knowledge Graph Question Answering (KGQA).
Knowledge Graphs (KG) is a set of factual triples, i.e., $\mathcal{G}=\{(h,r,t)\in\mathcal{V}\times\mathcal{R}\times\mathcal{V}\}$, where $h,r\in\mathcal{V}$ denote the head and tail entity, $r\in\mathcal{R}$ represents the relation.
Knowledge Graph Question Answering (KGQA) is a reasoning task that aims to predict answer entities $e_{a}\in\mathcal{A}_{q}$ based on $\mathcal{G}$. Following previous work Sun et al. (2019), we call entities mentioned in question $q$ as topic entities, denoted as $e_{t}\in\mathcal{T}_{q}$. Many datasets Talmor and Berant (2018); Yih et al. (2016b) give the standard SPARQL query of each question, which demonstrates a relation path from topic entity $e_{t}$ to answer entity $e_{a}$. We call this path as gold relation path, denote it as $w_{g}=e_{q}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}...\xrightarrow{r_{l}}e_{a}$. For example, the gold relation path of question in Figure 1 is $w_{g}=Apple\;Inc\xrightarrow{headquartered}Cupertino\xrightarrow{timezone}Pacific\;Standard\;Time$. In KGQA, $\forall i\in[1,l],\;(e_{i-1},r_{i},e_{i})\in\mathcal{G}$. That is, it is guaranteed that all triples in gold path are contained by $\mathcal{G}$.
4 Incomplete Knowledge Graph Question Answering (IKGQA)
4.1 Task Introduction
IKGQA differs from KGQA in that, in IKGQA, $\exists i\in[1,l],\;(e_{i-1},r_{i},e_{i})\notin\mathcal{G}$. That is, it doesn’t guarantee that all triples in gold path are contained by $\mathcal{G}$. For example, the triple (Cupertino, timezone, Pacific Standard Time) in $w_{g}$ may not be contained by $\mathcal{G}$. Therefore, models need to recall their inherent knowledge or reasoning from subgraph information.
4.2 Datasets Construction
We construct two IKGQA datasets based on two widely used KGQA datasets: WebQuestionSP (WebQSP) Yih et al. (2016b) and Complex WebQuestion (CWQ) Talmor and Berant (2018). Both datasets use Freebase Bollacker et al. (2008) as their background KG. To simulate incomplete KG, we delete some crucial triples, which appear in the gold relation path, of each question from the original KG. By doing this, simple semantic parsing methods almost fail to obtain the correct answers (except for some scenarios with multiple golden paths). In order to save computational cost, we selected the first 1,000 samples of these two datasets for constructing IKGQA questions.
For simplicity, we only consider one topic entity here. The process of generating crucial triples of a question is as follows:
  1. 1.
    Extract the topic entity $e_{q}$ from the question $q$ by LLMs.
  2. 2.
    Execute the SPARQL question $s_{q}$ (given in the datasets) on the background KG $\mathcal{G}$, and get all involved triples.
  3. 3.
    Filter property node (e.g., time, text, height), then convert the involved triples into a subgraph $g$.
  4. 4.
    Utilize breadth first search (BFS) to find the shortest path $p$, from topic entity $e_{q}$ to answer entity $a_{q}$.
  5. 5.
    Sample $k$ edges in the path $p$ as the crucial triples in the question.
5 Generate-on-Graph (GoG)
In this section, we introduce our method Generate-on-Graph (GoG), which can integrate the knowledge of KGs and LLMs, as well as utilize the reasoning ability of LLMs. The comparison between GoG and other previous methods is illustrated in Figure 3.
LLM as Agent. Motivated by ReAct Yao et al. (2023), we consider the LLM as an agent interacting with an environment to solve tasks. As shown in Figure 3 (d), for each step $i$, the agent first generates a thought $t_{i}\in\mathcal{L}$, where $\mathcal{L}$ is the language space, to decompose the original question (Thought 1) or decide which next sub-question should be solved (Thought 2). Then, based on the thought $t_{i}$, the agent generates an action $a_{i}\in\mathcal{A}$, where $\mathcal{A}$ is the action space, to search information by calling graph database API (Act 1, 2) or generate more information by reasoning and inherent knowledge (Act 4).
Action Space. GoG uses the selecting-generating-answering framework, which consists of three main actions: Search, Generate and Answer.
1. Search[target entity], which aims to find the most relevant top K neighbors of the target entity $e$. It is possible to search multiple entities at one time (Act 3 in Figure 3), for simplicity, we only consider searching one entity at one search action here. As there could be numerous neighbor entities of the target entity. We use the same strategy in ToG Sun et al. (2023) to filter unrelated neighbor entities. First, we retrieve all neighbor relations $r\in{R^{e}}$ linking to the target entity. This process is completed by pre-defined SPARQL queries. Then, we utilize the LLMs to select the top K relations that most related to the current sub-question (last thought). As shown in Act 2 of Figure 3 (d), given the target entity Cupertino, LLMs select the two relation {Located_in, Adjoin} from all neighbor relations. In the end, triples {(Cupertino, located_in, California), (Cupertino, adjoin, Palo Alto)} are appended to context as Obs 3.
2. Generate[sub-question], which make the agent generate new factual triples based on retrieval information and inherent knowledge. As shown in Act 4 of Figure 3 (d), although there is no triple directly representing the timezone of Cupertino, the agent can still infer the timezone of Cupertino based on {(Cupertino, located_in, California), (California, timezone, Pacific Standard Time)}. This process is similar to knowledge graph completion (KGC) Wang et al. (2017). It is also possible for the agent directly generates an entity, which is not explored before, based on its inherent knowledge. Therefore, we have to link the entity to its corresponding MID in the KG. This entity linking process is divided into two steps: (1) We retrieve some similar entities and their corresponding types based BM25 scores. (2) We utilize the LLM to select the most relevant entity based on the types.
3. Finish[answer], indicates that the agent finishes the task with answer. It should be noticed that the agent would also generate "Finish[unknown]", which means that there is not enough information for the agent to answer the question. In this case, we would roll back and search one more hop neighbors of the last target entity.
{tblr}
width = 0.8rows=rowsep=1pt, colspec = Q[452]Q[106]Q[125]Q[106]Q[125], row1 = c, row2 = c, row7 = c, column3 = c, column5 = c, cell12 = c=20.23, cell14 = c=20.23, cell21 = c=50.914, cell32 = c=20.23c, cell34 = c=20.23c, cell42 = c=20.23c, cell44 = c=20.23c, cell52 = c=20.23c, cell54 = c=20.23c, cell62 = c=20.23c, cell64 = c=20.23c, cell71 = c=50.914, cell82 = c, cell84 = c, cell92 = c, cell94 = c, cell102 = c, cell104 = c, cell112 = c,fg=MineShaft, cell113 = fg=MineShaft, cell114 = c,fg=MineShaft, cell115 = fg=MineShaft, cell122 = c, cell124 = c, cell132 = c, cell134 = c, cell142 = c, cell144 = c, cell152 = c, cell154 = c, cell162 = c, cell164 = c, cell172 = c, cell174 = c, vline2 = 1,3-6, vline4 = 1,3-6, vline2,4 = 8,10-11,13-17, hline1-3,7-10,12-13,18 = -, Method & CWQ WebQSP
without external knowledge
IO prompt 37.6 63.3
CoT 38.8 62.2
CoT+SC 45.4 61.1
RoG w/o planning 43.0 66.9
with external knowledge
CKG IKG-1 CKG IKG-1
w training
RoG 64.5 51.1 88.7 70.1
ChatKBQA 76.5 35.0 78.1 42.6
w/o training
KB-BINDER - - 50.7 34.4
StructGPT - - 76.4 53.7
ToG 47.2 36.7 76.9 60.6
GoG w/GPT-3.5 (Ours) 55.7 43.4 78.7 64.9
GoG w/GPT-4 (Ours) 75.2 61.0 84.4 74.8
Table 1: The Hits@1 scores of different models over two datasets in different settings (%). CKG and IKG-1 denote using complete and incomplete KG (dropping one crucial triple for each question), respectively. We use GPT-3.5 as backbone of KB-BINDER, StructGPT and ToG. Results of the other baselines were re-run by us, more details can be found in Appendix C. The boldface indicates the best result in the same setting.
6 Experiments
6.1 Experiments Setup
Evaluation Metrics
Following previous works Li et al. (2023d); Jiang et al. (2023); Sun et al. (2023), we use Hits@1 as our evaluation metric, which measures the proportion of questions whose top-1 predicted answer is correct.
Baselines
The baselines we compare can be divided into two groups: (1) LLM only methods, including standard prompting (IO prompt) Brown et al. (2020), Chain-of-Thought (CoT) prompting Wei et al. (2023) and Self-Consistency (SC) Wang et al. (2023b). (2) Semantic Parsing (SP) methods, including KB-BINDER Li et al. (2023c) and ChatKBQA Luo et al. (2023a). (3) Retrieval Augmented (RA) methods, including StructGPT Jiang et al. (2023), RoG Luo et al. (2023b) and ToG Sun et al. (2023), where RoG is the SOTA among all models requiring fine-tuning. All these methods are evaluated in both complete incomplete KGs.
Experiment Details
We use two LLMs as the backbone in our experiments: GPT-3.5 and GPT-4. We do not use Llama-2-70b-chat-hf Touvron et al. (2023), as we found that Llama’s output sometimes did not follow the format we defined in few-shot. We use OpenAI API to call GPT-3.5 and GPT-4. The maximum token length for each generation is set to 256. The temperature parameter is set to 0.7 in all experiments. We use 3 shots in GoG prompts for all the datasets. The prompts we use are listed in Appendix B. The details of re-running other baselines can be found in Appendix C.
{tblr}
width = colspec = Q[220]Q[180]Q[180]Q[180]Q[180], columneven = c, column3 = c, column5 = c, hlines, & IKG-1 IKG-2 IKG-3 IKG-4
CWQ 2.3 4.3 5.9 6.8
WebQSP 2.1 3.6 4.6 5.4
Table 2: The average number of edges deleted under different incompleteness degrees.
Datasets Details
For each dataset, we generate four incomplete KGs with varying degrees of completeness: IKG-1/2/3/4, representing randomly dropping 1/2/3/4 crucial triples in the KG. In addition to the crucial triples themselves, all relations between these two entities will also be deleted. The statistics of these incomplete KGs are shown in Table 2. Besides, we also ensure that after deleting these crucial triples, the number of neighbor nodes of the topic entities will not be zero, more details can be found in Appendix D.
      Method       WebQSP
      CKG       IKG-1       IKG-2       IKG-3       IKG-4
      StructGPT       76.4       52.5       49.1       47.9       46.4
      ToG       76.9       60.6       57.8       58.0       57.3
      GoG       78.7       62.2       60.9       59.6       57.4
      CWQ
      CKG       IKG-1       IKG-2       IKG-3       IKG-4
      ToG       47.2       36.7       33.1       32.2       31.7
      GoG       55.7       41.6       37.2       36.7       36.5
Table 3: The Hits@1 scores of prompt based methods (w/ GPT-3.5) under different numbers of missing triples (%). CKG represent using the complete KG. IKG-1/2/3/4 represent randomly dropping 1/2/3/4 crucial triples in the KG.
6.2 Main Results
Table 1 shows the Hits@1 scores of GoG and all baselines on two datasets in different settings. From the table, we can find that, compared with other prompt based methods, GoG can achieve the state-of-the-art performance on CWQ and WebQSP in both complete and incomplete KG settings.
In the CKG setting, the excellent performance of GoG mainly comes from its dynamic subgraph expansion strategy, which performs better than ToG in question involving compound value types (CVTs), an example of CVT is demonstrated in Figure 4. ToG is likely to think CVT nodes are not worthy to further explore and ignore them, as they do not offer information directly. Our GoG can easily solve this problem by expanding subgraph dynamically, that means if there is not enough information provided by the current subgraph, GoG would search one more hop, so the neighbors of CVT nodes is taken into consideration in this way.
In the IKG setting, the performance of SP methods, such as ChatKBQA and KB-BINDER, significantly declines. This is expected, as these SP methods don’t interact with the KGs, which means they have no idea of the absence of some triples. The performance of RoG, StructGPT and ToG also drop significantly. The performance of ToG and StructGPT on IKG is even worse than that without KG (IO prompt and CoT). That means, these methods still play a role of parsing to find answers rather than effectively integrating internal and external knowledge sources. Besides, in the IKG setting, these methods are even likely to retrieve wrong or unrelated paths which disturb LLMs. Our GoG can alleviate the problem by using the generate action, which utilizes the LLM to generate new factual triples when no direct answer is found after multiple rounds of searches on the KGs. Detailed analysis of answers generated by GoG can be checked in Appendix E.
Figure 4: An example of compound value types (CVTs) in Freebase dataset. Blue, green and orange nodes denote normal entities, CVT node and property node.
6.3 Performance under Different Numbers of Missing Triples
In order to explore the impact of different degrees of KG incompleteness on different methods, we evaluate the performance of methods (w/ GPT-3.5) under different numbers of missing triples, the results are demonstrated in Table 3.
It can be found that our GoG outperforms other prompt based methods consistently in different numbers of missing triples. Especially on the CWQ dataset, our GoG has a significant improvement on Hits@1 score, achieving average 7.5% improvement under all settings. That emphasizes the importance of integrate the external and inherent knowledge of LLMs. On the contrary, the performance of ToG on IKG is even much lower than that without KG, which indicates the performance of ToG still depends heavily on the completeness of KGs.
Furthermore, even though the majority of questions in the WebQSP dataset are single-hop questions, GoG can still demonstrate its advantages and perform better than ToG and StructGPT. This is because GoG can leverage the neighboring information of the topic entities to predict the tail entities while other methods can not make full use these information. More details can be found in Appendix F.
6.4 Performance with Different LLMs
Method WebQSP
CKG IKG-1 NKG
GoG w/GPT-3.5 78.7 64.9 62.2
GoG w/GPT-4 84.4 74.8 65.6
CWQ
CKG IKG-1 NKG
GoG w/GPT-3.5 55.7 43.4 38.8
GoG w/GPT-4 75.2 61.0 55.6
Table 4: The Hits@1 scores of GoG using different backbone models (%). CKG, IKG and NKG denote using complete, incomplete and no KG.
Figure 5: The Hits@1 scores of GoG with different generation strategies on the CWQ (a) and WebQSP (b) (%), w subgraph and w/o subgraph represent generating new factual triples with and without explored subgraph as context in the Generate action, respectively.
Figure 6: The Hits@1 scores of GoG with different number of related triples in the Generate action (%).
We evaluate how different backbone models affect GoG performance. Table 4 demonstrates that the performance of GoG using GPT-4 as backbone improves significantly. Especially under complete KGs setting, GoG (w/GPT-4) achieves 84.4 and 75.2 Hits@1 score on the WebQSP and CWQ datasets respectively, which achieve SOTA performance in prompt based methods and outperforms most fine-tuned methods.
In addition, we can also find that, the more powerful the backbone model, the better it utilizes the incomplete KGs. For example, on the WebQSP dataset, the performance of GoG (w/GPT-3.5) using incomplete KG is only 2.5% higher than that without using KG, while the improvement increases to 9.2% when using GPT-4 as backbone.
6.5 Ablation Study
The Effect of Explored Subgraphs
To explore the influence of explored subgraphs on the performance of GoG, we conduct experiments under different two Generate action (introduced in section 5) strategies: (1) Utilizing explored subgraphs, which is obtained in Search action, as context to generate new factual triples. (2) Generating new factual triples directly without explored subgraphs. As shown in Figure 5, GoG’s performance improves significantly with the help of explored subgraphs. This implies that, even incomplete KGs don’t provide answers directly, the related knowledge they provide is still helpful for models generating corresponding knowledge. There are two potential reasons: (1) Utilizing related subgraphs as context can activate LLMs’ memory corresponding to this knowledge. (2) LLMs can reason new factual triples based on the related subgraphs, as shown in Figure 3 (d).
The Effect of the Number of Related Triples
Aiming to find out how many related triples is required in the Generate action, we perform additional to find out how the number of related triples effect GoG’s performance. We select the most relevant k triples based on BM25. The results are shown in Figure 6. It can be observed that, GoG’s performance first increases and then decreases as the number of related triples increases. This is mainly because noisy and unrelated knowledge are introduced when the number of related triples is large. How to filter valuable knowledge from the explored subgraph could be a future direction.
7 Conclusion
In this paper, we propose leveraging LLMs for QA under Incomplete Knowledge Graph (IKGQA), which is closer to real-world scenarios, and construct relevant datasets. We propose Generate-on-Graph (GoG), which can effectively integrate the external and inherent knowledge of LLMs. Experiments on two datasets show the superiority of GoG, and demonstrate that an incomplete KG can still help LLMs answer complex questions.
Limitation
The limitations of our proposed GoG are as follows: (1) We only evaluate GoG on CWQ and WebQSP, which uses Freebase as their background KG. The experimental data may comes from some specific domains. (2) It is possible for LLM to hallucinate in the generation process, which is unavoidable for existing LLMs. (3) There is room for further improvement in performance, as GoG’s performance is lower than that with CoT prompt when KGs are very incomplete.
Ethics Statement
This paper proposes a method for complex question answering in incomplete knowledge graph, and the experiments are conducted on public available datasets. As a result, there is no data privacy concern. Meanwhile, this paper does not involve human annotations, and there are no related ethical concerns.
Appendix A Semantic Parsing Methods Details
The training datasets for SP methods are constructed under the complete KGs, which means that "Time Zone" corresponds directly to the relation "ns:location.location.time_zones" rather than a two-hop path "ns:location.located_in -> ns:location.location.time_zones". An example in CWQ is shown in Table 6. This means SP models trained on CWQ will always output "?c ns:location.location.time_zones ?x" instead of "?c ns:location.located_in ?y . ?y ns:location.location.time_zones ?x". Therefore, these methods will fail under Incomplete KGs. In another word, semantic parsing methods don’t interact with the KGs, which means they have no idea of the absence of some triples.
{tblr}
width = 0.9colspec = Q[104]Q[837], cell12 = fg=MineShaft, cell21 = fg=MineShaft, vline2 = -, hline1-3 = -, Question & In the nation that spends the Bahamian dollar as currency, what time zone is used?
SPARQL PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT DISTINCT ?x WHERE {
  FILTER (?x != ?c) FILTER (!isLiteral(?x) OR lang(?x) = ” OR langMatches(lang(?x), ’en’))
  ?c ns:location.country.currency_used ns:m.01l6dm .
  ?c ns:location.location.time_zones ?x .
Table 6: An example about "timezone" in the CWQ dataset.
Appendix B Prompt List
The prompts used in GoG are shown in Table 7.
{tblr}
width = colspec = Q[90]Q[817]Q[27]Q[27], cell11 = c, cell12 = c, hline1-5 = 1-2, Tasks & Prompt
GoG Instruction Solve a question answering task with interleaving Thought, Action, Observation steps. Thought can reason about the current situation, and Action can be three types:
(1) Search[entity1 | entity2 | …], which searches the exact entities on Freebase and returns their one-hop subgraphs. You should extract the all concrete entities appeared in your last thought without redundant words, and you should always select entities from topic entities in the first search.
(2) Generate[thought], which generate some new triples related to your last thought. These new triples may come from your inherent knowledge directly or reasoning from the given triples.
(3) Finish[answer1 | answer2 | …], which returns the answer and finishes the task. The answers should be complete entity label appeared in the triples. If you don’t know the answer, please output Finish[unknown].
Entities and answers should be separated by tab.
Attention please, entities begin with "m." (e.g., m.01041p3) represent CVT (compound value type) node, and they shouldn’t be selected as the final answers. To find out those entities involved in these event, you could select them as the entities to be searched.
You should generate each step without redundant words.
Here are some examples.
In-Context Few-shot
Question: {Question}
Topic Entity: {List of Topic Entities}
Thought 1:
Filter Relations Please select 3 relations that most relevant to the question and rank them. You should answer these relations in list format directly without redundant words.
Here are some examples.
In-Context Few-shot
Thought: {Thought}
Entity: {Entity}
Relation: {List of Relations}
Answer:
Generate Triples Given the existing triples, please generate some new triples related to your current thought. These new triples may come from your inherent knowledge directly or reasoning from the given triples.
Here are some examples.
In-Context Few-shot
Thought: {Thought}
Known Triples: {Explored Triples}
Generated Triples:
Table 7: Prompts for different tasks used in GoG.
Appendix C Settings for Baselines
Following ToG, the Freebase dump is acquired from https://developers.google.com/freebase?hl=en, we deploy Freebase with Virtuoso. GoG, RoG, KB-BINDER and ChatKBQA are evaluated on the same Freebase database.
RoG. We use the checkpoints and the default settings provided by the official repository: n_beam=3 in generating rule, max_new_tokens=512 in inferring answers.
ChatKBQA. We use the predicted S-expression provided by the official repository, and convert them into SPARQL queries. To compare ChatKBQA with other models fairly, we execute these SPARQL queries under the Freebase database mention before instead the DB files provided by them. Therefore, the performance of ChatKBQA reported in Table 1 is slightly different from that in their original paper.
KB-BINDER. We use the official repository and use KB-BINDER (6)-R (with majority vote and retrieve the most similar exemplars) to infer answers. However, the code-davinci-002 used in their original paper is not available, so we use GPT-3.5 instead. Besides, to reduce runtime, we decreased the number of candidate MID combinations (despite that, it still takes about 4 hours to answer 200 questions). Therefore, the performance of KB-BINDER reported in Table 1 is slightly different from that in their original paper.
ToG. We use the official repository and their default settings for inferring answers: max_length=256, width=3, depth=3. Since the official repository doesn’t provide the alias answers in the CWQ dataset, we evaluate ToG on the CWQ dataset without considering alias answers (the same strategy for all models). Therefore, the performance of ToG reported in Table 1 is slightly different from that in their original paper.
StructGPT. We use the official repository and running scripts to evaluate StructGPT on the WebQSP dataset.
Appendix D Statistics of Topic Entities in IKGs
The statistics of topic entities are shown in Table 8, and we drop those samples which have isolated topic entities (topic entity without any neighbor node).
{tblr}
width = 0.9colspec = Q[108]Q[400]Q[80]Q[80]Q[80]Q[80], column3 = c, column4 = c, column5 = c, column6 = c, cell21 = r=2, cell22 = fg=MineShaft, cell41 = r=2, cell42 = fg=MineShaft, hline1,2,4,6 = -, Dataset & IKG-1 IKG-2 IKG-3 IKG-4
CWQ Median number of neighbor nodes 27 26 27 27
Number of isolated topic entities 19 42 59 53
WebQSP Median number of neighbor nodes 428 427 427 426
Number of isolated topic entities 1 2 1 2
Table 8: Statistics of topic nodes in Incomplete KGs. Isolated topic entity represent topic entity without any neighbor node.
Appendix E Result Analysis
Table 9 illustrates the frequency of the Generate operation in different datasets alongside their corresponding Hits@1 scores. In the complete KGs setting, GoG still conducts the Generate operation when related relations are not correctly selected or when answers to sub-questions cannot be directly found via a one-hop relationship. In the incomplete KGs setting, the frequency of the Generate operation is higher, as GoG needs to generate new factual triples that are missing in the KGs. Hits@1 scores under both settings mean that most generation leading to correct results.
We considered four types of errors: (1) Knowledge Missing. (2) Hallucination. (3) Error before Generate. (4) False Negative. The distribution is shown in Figure 7. It can be found that there are indeed about 26% error are caused by hallucination. Besides, the Figure 6 demonstrates that the performance of the model improves with the number of relevant triples introduced in the Generation step, which means introducing more related triples can alleviate hallucination problems to a certain extent.
Figure 7: The error proportions of GoG under IKG-1 of CWQ.
{tblr}
width = 0.9colspec = Q[185]Q[181]Q[271]Q[271], cells = c, cell13 = fg=MineShaft, cell14 = fg=MineShaft, cell21 = r=2, cell23 = fg=MineShaft, cell24 = fg=MineShaft, cell33 = fg=MineShaft, cell41 = r=2, vline2-4 = 1-5, vline3-4 = 3,5, hline1-2,4,6 = -, Dataset & Models CWQ WebQSP
CKG GPT-3.5 17.7% (55.9) 17.1% (66.6)
GPT-4 4.6% (54.3) 9.7% (65.9)
IKG-1 GPT-3.5 23.7% (57.3) 24.1% (63.4)
GPT-4 8.8% (57.9) 18.4% (70.6)
Table 9: Times of Generate operation in different KG settings. Numbers in brackets represent corresponding Hits@1 score.
Appendix F Case Study
In this section, we present a case analysis to evaluate the utility of GoG, as demonstrated in Table 10. In this case, GoG will first search the neighbors of "Appalachian Mountains" and obtain <Bald Eagle Mountain, mountain.mountain_range, Appalachian Mountains>, <Spaulding Mountain, mountain.mountain_range, Appalachian Mountains> and <Old Rag Mountain, mountain.mountain_range, Appalachian Mountains>, then GoG can infer that Appalachian Mountains are also located in North America simply by knowing information about the other three mountains. However, ToG fails on this question once the crucial triple <Appalachian Mountains, location.location.containedby, North America> is missing. In another word, for one-hop questions, GoG can use the neighbors information of the topic entities to predict the tail entities while other methods can not make full use of the neighbors information.
{tblr}
width = colspec = Q[194]Q[748], cell22 = fg=MineShaft, hlines, vline2 = -, Question & Where do the appalachian mountains run?
Dropped Crucial Triple  Appalachian Mountains, location.location.containedby, North America
ToG Reasoning Chains: None
Prediction: I’m sorry, but I need more information to answer your question.
GoG Thought 1: I need to search the Appalachian Mountains to find out where they run.
Act 1: Search[Appalachian Mountains]
Obs 1: Appalachian Mountains, location.contains, Brasstown Bald
Appalachian Mountains, location.contains, Craggy Dome
Appalachian Mountains, location.geolocation, m.02_qbv6
Bald Eagle Mountain, mountain.mountain_range, Appalachian Mountains
Spaulding Mountain, mountain.mountain_range, Appalachian Mountains
Old Rag Mountain, mountain.mountain_range, Appalachian Mountains
Thought 2: The Appalachian Mountains contain various locations and are part of geolocations. To find out where they run, I need to generate some new triples based on these given triples and my inherent knowledge.
Act 2:  Generate[where do the Appalachian Mountains run]
Obs 2: Appalachian Mountains, geographical_feature.contains_location, Blue Ridge Mountains
Appalachian Mountains, geographical_feature.start_location, Alabama
Appalachian Mountains, geographical_feature.start_location, North America
Thought 3: The Appalachian Mountains start from North America
Act 3: Finish[Alabama | North America ]
Ground Truth North America
Table 10: Comparison between ToG and GoG.